OpenStudy (anonymous):

is 2cos^2-1 equal to 1 or is it cos^2-1 equal to 1???

OpenStudy (amistre64):

neither?

OpenStudy (amistre64):

2cos^2 - 1 is more apt to equal 1 in the long run; but as a rule; no

OpenStudy (turingtest):

\[1-\cos^2 x=\sin^2 x\]

OpenStudy (anonymous):

not if x = 0

OpenStudy (anonymous):

kk, lim x->0 tan2x/x

OpenStudy (anonymous):

tan2x/x=sin2x/cos2x/x correct?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

well, maybe if i read it right

OpenStudy (anonymous):

one sec ill draw it..

OpenStudy (amistre64):

2sin(x)cos(x) ------------------- xcos^2(x) -xsin^2(x) maybe

OpenStudy (amistre64):

sin(x)/x limits to ... 1 or 0 i cnat recall

OpenStudy (anonymous):